OpenStudy (anonymous):

Evaluate the expression. 10 P 4 A. 5040 B. 720 C. 1 D. 10

OpenStudy (cwrw238):

= 10*9*8*7

OpenStudy (cwrw238):

=90 * 56

OpenStudy (campbell_st):

use the formula \[^{n}P _{k} = \frac{n!}{(n - k)!}\] n = 10 and k = 4

OpenStudy (anonymous):

so 5040

OpenStudy (cwrw238):

right

OpenStudy (cwrw238):

i use a 'short cut' to do these but you should know formula

OpenStudy (cwrw238):

10 P 3 for example = 10*9*8

OpenStudy (anonymous):

i understand thank you so a 11 p 5 is 11*10*9*8*7

OpenStudy (cwrw238):

yea and also 11 P 8 is same as 11 P 3 (11-8 = 3) - makes that easier to work out

OpenStudy (cwrw238):

for combinations where the order doesn't matter you proceed like this 10 C 4 = 10*9*8*7 ------- 4*3*2*1 - number of digits on bottom = number on top